OpenStudy (anonymous):

Solve for x. 6 – 3(2x – 4) = 2(1 + x) A. –2 B. 0 C. –1 D. 2

OpenStudy (anonymous):

Use the distributive law. and balance each side of the equation.

OpenStudy (yanasidlinskiy):

Distribute the other side! 2+2x

OpenStudy (anonymous):

6 – 3(2x – 4) = 2(1 + x) Expand the brackets to get: 6 - 6x +12 = 2 + 2x Collect like terms 6 + 12 - 2 = 2x + 6x 16 = 8x div both sides by 8 x = 2

OpenStudy (anonymous):

thnx yall